# Prompt-CAM: A Simpler Interpretable Transformer for Fine-Grained Analysis

### v1

@@ Abstract
We present a simple usage of pre-trained Vision Transformers (ViTs) for fine-grained analysis, aiming to identify and localize the traits that distinguish visually similar categories, such as different bird species or dog breeds. Pre-trained ViTs such as DINO have shown remarkable capabilities to extract localized, informative features. However, using saliency maps like Grad-CAM can hardly point out the traits: they often locate the whole object by a blurred, coarse heatmap, not traits. We propose a novel approach Prompt Class Attention Map (Prompt-CAM) to the rescue. Prompt-CAM learns class-specific prompts to a pre-trained ViT and uses the corresponding outputs for classification. To classify an image correctly, the true-class prompt must attend to the unique image patches not seen in other classes’ images, i.e., traits. As such, the true class’s multi-head attention maps reveal traits and their locations. Implementation-wise, Prompt-CAM is almost a free lunch by simply modifying the prediction head of Visual Prompt Tuning (VPT). This makes Prompt-CAM fairly easy to train and apply, sharply contrasting other interpretable methods that design specific models and training processes. It is even simpler than the recently published INterpretable TRansformer (INTR), whose encoder-decoder architecture prevents it from leveraging pre-trained ViTs. Extensive empirical studies on a dozen datasets from various domains (e.g., birds, fishes, insects, fungi, flowers, food, and cars) validate Prompt-CAM superior interpretation capability.
@@ 1 Introduction
Figure 1: Prompt-CAM identifies and localizes traits class-specifically and consistently. (a) Images of “Baltimore Oriole.” (b) Top-2 “Baltimore Oriole”-specific attention maps (across 12 heads) using Prompt-CAM, focusing on the orange belly and black wings. (c) Top-2 “Bronzed Cowbird”-specific attention maps on the same images, focusing on the black wings and heads.
What if we look at the attention maps? ViTs rely on self-attention to relate image patches; the [CLS] token aggregates image features by attending to informative patches. As shown in [7, 38, 26], the attention maps of the [CLS] token do highlight local regions inside the object. However, these regions are not “class-specific.” Instead, they often point to the same object regions across different categories, such as body parts like heads, wings, and tails of bird species. While these are where traits usually reside, they are not traits. For example, the distinction between “Red-winged Blackbird” and other bird species is the red spot on the wing, having little to do with other body parts.
@@
Our proposal is to prompt the ViTs with learnable “class-specific” tokens, one for each class, inspired by [30, 19, 48]. These “class-specific” tokens, once inputted to ViTs, attend to image patches via self-attention, just like the [CLS] token. Yet, unlike [CLS] token which is “class-agnostic,” these “class-specific” tokens can attend to the same image differently, having the potential to highlight local regions that are specific to the corresponding classes, i.e., traits.
@@
We implement our approach, which we name Prompt Class Attention Map (Prompt-CAM), as follows. Given a pre-trained ViT and a fine-grained classification dataset with $C$ classes, we add $C$ learnable tokens as extra inputs to the first Transformer layer. To make these tokens “class-specific,” we collect their corresponding output vectors after the final Transformer layer and perform inner products with a shared vector (also learnable) to obtain $C$ “class-specific” scores, following [30]. One may interpret each class-specific score as how clearly the corresponding class’s traits are seen in the input image. Intuitively, the input image’s ground-truth class should possess the highest score, and we encourage this by minimizing a cross-entropy loss, treating the scores as logits. We keep the whole pre-trained ViT frozen and only optimize the $C$ tokens and the shared scoring vector. See section 3 for details and variants.
@@
For interpretation during inference, we input the image and the $C$ tokens simultaneously to the ViT to obtain the $C$ scores. One can then select a specific class, for example, the highest-score class, and visualize the $R$ multi-head attention maps to the image patches. (See section 3 for how to rank these $R$ maps to highlight the most discriminative traits.) When the highest-score class is the ground-truth class, the attention maps reveal its traits. Otherwise, comparing the highest-score class’s attention and the ground-truth class’s attention explains why the image is misclassified. Example reasons involve the object in the image being partially occluded or in an odd pose, such that its traits were invisible, or somehow its appearance being too similar to a wrong class, perhaps due to lighting conditions (see Figure 5).
@@
Prompt-CAM is fairly easy to implement and train. It requires no change to pre-trained ViTs and no specially designed loss function or training strategy—just the standard cross-entropy loss and SGD. Indeed, building upon Visual Prompt Tuning (VPT) [12], one merely needs to adjust a few lines of code and can enjoy fine-grained interpretation. This simplicity sharply contrasts other interpretable methods like ProtoPNet [6] and ProtoTree [25]. Compared to INterpretable TRansformer (INTR) [30], which also featured simplicity, Prompt-CAM has three notable advantages. First, Prompt-CAM is encoder-only and can potentially take any ViT encoders. In contrast, INTR is built upon an encoder-decoder model pre-trained on object detection datasets. Thus, Prompt-CAM can leverage up-to-date pre-trained models more easily. Second, as a result, Prompt-CAM can be trained much faster—only the prompts and the shared vector need to be learned. In contrast, INTR typically demands full fine-tuning. Third, Prompt-CAM produces cleaner and sharper attention maps than INTR, which we attribute to the usages of state-of-the-art ViTs like DINO or DINOv2. Put things together, we view Prompt-CAM as a simpler yet stronger interpretable Transformer.
(a)                                    (b)
@@
We validate Prompt-CAM on over a dozen datasets: CUB-200-2011 [43], Birds-525 [32], Oxford Pet [29], Stanford Dogs [15], Stanford Cars [16], iNaturalist-2021-Moths [41], Fish Vista [23], Rare Species [40], Insects-2 [47], iNaturalist-2021-Fungi [41], Oxford Flowers [27], Medicinal Leaf [35], Stanford Cars [16] and Food 101 [2]. Prompt-CAM can identify different traits of a category through multi-head attention and consistently localize them in images. To our knowledge, Prompt-CAM is the only explainable or interpretable method for vision that has been evaluated on such a broad range of domains. We further show Prompt-CAM’s extendability by applying it to discovering taxonomy keys. Our contributions are two-fold.
@@
2 Related Work (See the Suppl. for Details)
Seeing the remarkable performance of neural networks, researchers have strived to understand their predictions. Explainable methods do not change the neural network models but derive saliency maps or image masks to pinpoint where the model looks [36, 13, 24, 44]. Interpretable methods design specific neural networks (and training procedures) whose inner workings reveal how the model makes predicts [45, 30, 20, 49]. Prompt-CAM belongs to the interpretable methods but makes minimal changes to ViTs and can be easily trained and applied.
@@ 3.2 Prompt Class Attention Map (Prompt-CAM)
Prompt-CAM-Deep. While straightforward, Prompt-CAM-Shallow has two potential drawbacks. First, the class-specific prompts attend to every layer’s patch features, i.e., $\bm{E}_{i}$, $i=0,\cdots,N-1$. However, features of the early layers are often not informative enough but noisy for differentiating classes. Second, the prompts $\bm{p}^{1},\cdots,\bm{p}^{C}$ have a “double duty.” Individually, each needs to highlight class-specific traits. Collectively, they need to adapt pre-trained ViTs to downstream tasks, the original purpose of VPT [12]. In our case, the downstream task is a new usage of ViTs on a specific fine-grained dataset.
@@ 3.3 Trait Identification and Localization
Which traits are more discriminative? For categories that are so distinctive like “Red-winged Blackbird,” a few traits are sufficient to distinguish them from others. To automatically identify these most discriminative traits, we take a greedy approach, progressively blurring the least important traits until the image is classified wrongly. The remaining ones highlight the traits that are sufficient for classification.
Suppose class $c$ is the true class and the image is correctly classified. In each greedy step, for each of the unblurred heads indexed by $r$, we iteratively replace ${\bm{\alpha}}^{c,r}_{N-1}$ with $\frac{1}{M}\textbf{1}$ and recalculate $s[c]$ in Equation 1, where $\textbf{1}\in\mathbb{R}^{M}$ is an all-one vector. Doing so essentially blurs the $r$-th head for class $c$, preventing it from focusing. The head with the highest blurred $s[c]$ is thus the least important, as blurring it hurts classification the least. See Suppl. for more details.
@@ 3.4 Variants and Extensions
Focused Prompt-CAM for pair-wise comparison. To claim the highest prediction score, the ground-truth class’s prompt $\bm{p}_{N-1}^{y}$ must identify a comprehensive set of traits to tell itself apart from all the other categories. As a result, it may not prioritize the nuanced differences in patterns, colors, and shapes that distinguish it from the most visually similar class (e.g., animal species under the same Genus).
We investigate one approach to mitigating this, which is to feed only the ground-truth and targeted reference classes’ prompts into Prompt-CAM-Deep during inference, denoted by $y$ and $y^{\prime}$, respectively. We note that the MSA block in the Transformer layer $L_{N}$ allows each class-specific prompt to attend to the others. Subsampling the input prompts thus wound change the output score $s[y]$ in Equation 1 and consequently the order of the least important heads, focusing the model on what to look at.
@@ 3.5 What is Prompt-CAM suited for?
Prompt-CAM is designed with such a mindset. Unlike conventional classifiers with a fully connected layer on top, Prompt-CAM follows [30] to learn a shared vector $\bm{w}$ in Equation 1, whose goal is NOT to record class-specific information BUT to answer the “binary” question: Based on where a class-specific prompt attends, does the class find itself in the input image?
@@
where $\bm{w}_{c}$ stores the fully connected weights for class $c$. We argue that such a formulation offers a “detour” such that the model can correctly classify an image $\bm{I}$ of class $y$ even without meaningful attention weights. In essence, the model can choose to produce holistically discriminative value features from $\bm{I}$ with no spatial resolution, such that $\bm{v}^{j}$ aligns with $\bm{w}_{y}$ but $\bm{v}^{j}=\bm{v}^{j^{\prime}},\forall j\neq j^{\prime}$. In this case, no matter what ${\bm{\alpha}}^{\star}$ is, as long as it sums to one as default in the $\operatorname{softmax}$ formula, the prediction remains intact.
@@
  • Generate representative, localized features $\bm{v}^{j}$ to precisely encode the trait (e.g., red spots) within each patch $j$.
  • Generate distinctive attention weights ${\bm{\alpha}}^{c}$ among classes; each highlights traits frequently seen in class $c$.
@@
Figure 4: Visualization of Prompt-CAM on eight datasets. We show the top four attention maps (from left to right) per correctly classified test example triggered by the ground-truth classes. As top head indices per image may vary, traits may not align across rows.
@@ 4.1 Experimental Setup
Model. We consider three pre-trained ViT backbones, DINO [4], DINOv2 [28], and BioCLIP [37] across different scales including ViT-B (the main one we use) and ViT-S. The backbones are kept completely frozen when applying Prompt-CAM. More details can be found in Suppl.
@@ 4.2 Experiment Results
Is Prompt-CAM faithful? For the faithfulness analysis, we apply Grad-CAM [36], Eigen-CAM [24], and Layer-CAM [13] to generate post-hoc saliency maps on the Linear Probing classifier as baselines. We then compare these with Prompt-CAM using the insertion and deletion metrics [31]. As shown in Table 1, Prompt-CAM yields significantly higher insertion scores and lower deletion scores, indicating a stronger focus on discriminative image traits and highlighting Prompt-CAM’s enhanced interpretability over standard post-hoc algorithms.
@@
Grad-CAM 0.52 0.16
Layer-CAM 0.54 0.13
Eigen-CAM 0.42 0.33
@@
Prompt-CAM excels in trait identification (human assessment). Human assessment is conducted to evaluate trait identification quality for Prompt-CAM, TesNet [45], and ProtoConcepts [20]. Participants with no prior knowledge about the algorithms were instructed to compare the expert-identified traits (in text, such as orange belly) and the top heatmaps generated by each method. If an expert-identified trait is seen in the heatmaps, it is considered identified by the algorithm. On average, participants recognized $60.49\%$ of traits for Prompt-CAM, significantly outperforming TesNet and ProtoConcepts whose recognition rates are $39.14\%$ and $30.39$%, respectively. The results highlight Prompt-CAM’s superiority in emphasizing and conveying relevant traits effectively. More details are in Suppl.
Classification accuracy comparison. Building on the discussion in subsection 3.5, we observe that Prompt-CAM shows a slight accuracy drop compared to Linear Probing (see Table 2). However, the images misclassified by Prompt-CAM but correctly classified by Linear Probing align with our design philosophy: Prompt-CAM classifies images based on the presence of class-specific, localized traits and would fail if they are invisible. As illustrated in Figure 5, discriminative traits—such as the red breast of the Red-breasted Grosbeak—are barely visible in images misclassified by Prompt-CAM due to occlusion, unusual poses, or lighting conditions. Linear Probing correctly classifies them by leveraging global information such as body shapes and backgrounds. Please see more analysis in Suppl.
Figure 5: Images misclassified by Prompt-CAM but correctly classified by Linear Probing. Species-specific traits—such as the red breast of Red-breasted Grosbeak—are barely visible in misclassified images while Linear Probing uses global features such as body shapes, poses, and backgrounds for correct predictions.
@@
Linear Probing 98.1 78.6 82.4 92.4
Prompt-CAM 97.4 71.9 77.0 87.6
@@
Figure 6: Comparison of interpretable models. Visual demonstration (heatmaps and bounding boxes) of the four most activated responses of attention heads (Prompt-CAM and INTR) or prototypes of each method on a Lazuli Bunting example image.

### v2

# Prompt-CAM: Making Vision Transformers Interpretable for Fine-Grained Analysis

@@ Abstract
We present a simple approach to make pre-trained Vision Transformers (ViTs) interpretable for fine-grained analysis, aiming to identify and localize the traits that distinguish visually similar categories, such as bird species. Pre-trained ViTs, such as DINO, have demonstrated remarkable capabilities in extracting localized, discriminative features. However, saliency maps like Grad-CAM often fail to identify these traits, producing blurred, coarse heatmaps that highlight entire objects instead. We propose a novel approach, Prompt Class Attention Map (Prompt-CAM), to address this limitation. Prompt-CAM learns class-specific prompts for a pre-trained ViT and uses the corresponding outputs for classification. To correctly classify an image, the true-class prompt must attend to unique image patches not present in other classes’ images (i.e., traits). As a result, the true class’s multi-head attention maps reveal traits and their locations. Implementation-wise, Prompt-CAM is almost a “free lunch,” requiring only a modification to the prediction head of Visual Prompt Tuning (VPT). This makes Prompt-CAM easy to train and apply, in stark contrast to other interpretable methods that require designing specific models and training processes. Extensive empirical studies on a dozen datasets from various domains (e.g., birds, fishes, insects, fungi, flowers, food, and cars) validate the superior interpretation capability of Prompt-CAM. The source code and demo are available at https://github.com/Imageomics/Prompt_CAM.
Figure 1: Illustration of Prompt-CAM. By learning class-specific prompts for a pre-trained Vision Transformer (ViT), Prompt-CAM enables multiple functionalities. (a) Prompt-CAM achieves fine-grained image classification using the output logits from the class-specific prompts. (b) Prompt-CAM enables trait localization by visualizing the multi-head attention maps queried by the true-class prompt. (c) Prompt-CAM identifies common traits shared between species by visualizing the attention maps queried by another-class prompt. (d) Prompt-CAM can identify the most discriminative traits per species (e.g., distinctive yellow chest and black neck for “Scott Oriole”) by systematically masking out the least important attention heads. See subsection 2.3 for details.
@@ 1 Introduction
What if we look at the attention maps? ViTs rely on self-attention to relate image patches; the [CLS] token aggregates image features by attending to informative patches. As shown in [7, 39, 27], the attention maps of the [CLS] token do highlight local regions inside the object. However, these regions are not “class-specific.” Instead, they often focus on the same object regions across different categories, such as body parts like heads, wings, and tails of bird species. While these are where traits usually reside, they are not traits. For example, the distinction between “Red-winged Blackbird” and other bird species is the red spot on the wing, having little to do with other body parts.
@@
Our proposal is to prompt ViTs with learnable “class-specific” tokens, one for each class, inspired by [31, 20, 50, 17]. These “class-specific” tokens, once inputted into ViTs, attend to image patches via self-attention, similar to the [CLS] token. However, unlike the [CLS] token, which is “class-agnostic,” these “class-specific” tokens can attend to the same image differently, with the potential to highlight regions specific to the corresponding classes, i.e., traits.
@@
We implement our approach, named Prompt Class Attention Map (Prompt-CAM), as follows. Given a pre-trained ViT and a fine-grained classification dataset with $C$ classes, we add $C$ learnable tokens as additional inputs alongside the input image. To make these tokens “class-specific,” we collect their corresponding output vectors after the final Transformer layer and perform inner products with a shared vector (also learnable) to obtain $C$ “class-specific” scores, following [31]. One may interpret each class-specific score as how clearly the corresponding class’s traits are visible in the input image. Intuitively, the input image’s ground-truth class should possess the highest score, and we encourage this by minimizing a cross-entropy loss, treating the scores as logits. We keep the whole pre-trained ViT frozen and only optimize the $C$ tokens and the shared scoring vector. See section 2 for details and variants.
@@
For interpretation during inference, we input the image and the $C$ tokens simultaneously to the ViT to obtain the $C$ scores. One can then select a specific class (e.g., the highest-score class) and visualize its multi-head attention maps over the image patches. See Figure 1 for an illustration and section 2 for how to rank these maps to highlight the most discriminative traits. When the highest-score class is the ground-truth class, the attention maps reveal its traits. Otherwise, comparing the attention maps of the highest-score class with those of the ground-truth class helps explain why the image is misclassified. Possible reasons include the object being partially occluded or in an unusual pose, making its traits invisible, or the appearance being too similar to a wrong class, possibly due to lighting conditions (Figure 5).
@@
Prompt-CAM is fairly easy to implement and train. It requires no change to pre-trained ViTs and no specially designed loss function or training strategy—just the standard cross-entropy loss and SGD. Indeed, building upon Visual Prompt Tuning (VPT) [12], one merely needs to adjust a few lines of code and can enjoy fine-grained interpretation. This simplicity sharply contrasts other interpretable methods like ProtoPNet [6] and ProtoTree [26]. Compared to INterpretable TRansformer (INTR) [31], which also featured simplicity, Prompt-CAM has three notable advantages. First, Prompt-CAM is encoder-only and can potentially utilize any ViT encoder. In contrast, INTR is built upon an encoder-decoder model pre-trained on object detection datasets. As a result, Prompt-CAM can more easily leverage up-to-date pre-trained models. Second, Prompt-CAM can be trained much faster—only the prompts and the shared vector need to be learned. In contrast, INTR typically requires full fine-tuning. Third, Prompt-CAM produces cleaner and sharper attention maps than INTR, which we attribute to the use of state-of-the-art ViTs like DINO [4] or DINOv2 [29]. Taken together, we view Prompt-CAM as a simpler yet more powerful interpretable Transformer.
@@
We validate Prompt-CAM on over a dozen datasets: CUB-200-2011 [45], Birds-525 [33], Oxford Pet [30], Stanford Dogs [15], Stanford Cars [16], iNaturalist-2021-Moths [43], Fish Vista [24], Rare Species [41], Insects-2 [49], iNaturalist-2021-Fungi [43], Oxford Flowers [28], Medicinal Leaf [36], Stanford Cars [16], and Food 101 [2]. Prompt-CAM can identify different traits of a category through multi-head attention and consistently localize them in images. To our knowledge, Prompt-CAM is the only explainable or interpretable method for vision that has been evaluated on such a broad range of domains. We further show Prompt-CAM’s extendability by applying it to discovering taxonomy keys. Our contributions are two-fold.
@@
Due to space constraints, we provide a detailed related work section in the Supplementary Material (Suppl.).
@@ 2.2 Prompt Class Attention Map (Prompt-CAM)
Prompt-CAM-Deep. While straightforward, Prompt-CAM-Shallow has two potential drawbacks. First, the class-specific prompts attend to every layer’s patch features, i.e., $\bm{E}_{i}$, $i=0,\cdots,N-1$. However, features of the early layers are often not informative enough but noisy for differentiating classes. Second, the prompts $\bm{p}^{1},\cdots,\bm{p}^{C}$ have a “double duty.” Individually, each needs to highlight class-specific traits. Collectively, they need to adapt pre-trained ViTs to downstream tasks, which is the original purpose of VPT [12]. In our case, the downstream task is a new usage of ViTs on a specific fine-grained dataset.
@@ 2.3 Trait Identification and Localization
Which traits are more discriminative? For categories that are so distinctive, like “Red-winged Blackbird,” a few traits are sufficient to distinguish them from others. To automatically identify these most discriminative traits, we take a greedy approach, progressively blurring the least important attention maps until the image is misclassified. The remaining ones highlight traits that are sufficient for classification.
Suppose class $c$ is the true class and the image is correctly classified. In each greedy step, for each of the unblurred heads indexed by $r^{\prime}$, we iteratively replace ${\bm{\alpha}}^{c,r^{\prime}}_{N-1}$ with $\frac{1}{M}\textbf{1}$ and recalculate $s[c]$ in Equation 1, where $\textbf{1}\in\mathbb{R}^{M}$ is an all-one vector. Doing so essentially blurs the $r^{\prime}$-th head for class $c$, preventing it from focusing. The head with the highest blurred $s[c]$ is thus the least important, as blurring it degrades classification the least. See Suppl. for details.
@@ 2.5 What is Prompt-CAM suited for?
Prompt-CAM is designed with this mindset. Unlike conventional classifiers that employ a fully connected layer on top, Prompt-CAM follows [31] and learns a shared vector $\bm{w}$ in Equation 1. The goal of $\bm{w}$ is NOT to capture class-specific information BUT to answer a “binary” question: Based on where a class-specific prompt attends, does the class recognize itself in the input image?
@@
where $\bm{w}_{c}$ stores the fully connected weights for class $c$. We argue that this formulation allows for a potential “detour,” enabling the model to correctly classify an image $\bm{I}$ of class $y$ even without meaningful attention weights. In essence, the model can choose to produce holistically discriminative value features from $\bm{I}$ without preserving spatial resolution, such that $\bm{v}^{j}$ aligns with $\bm{w}_{y}$ but $\bm{v}^{j}=\bm{v}^{j^{\prime}},\forall j\neq j^{\prime}$. In this case, regardless of the specific values of ${\bm{\alpha}}^{\star}$, as long as they sum to one—as is default in the $\operatorname{softmax}$ formulation—the prediction remains unaffected.
@@
  • Generate localized features $\bm{v}^{j}$ that highlight discriminative patches (e.g., the red spot on the wing) of an image.
  • Generate distinctive attention weights ${\bm{\alpha}}^{c}$ across classes, each focusing on traits frequently seen in class $c$.
@@
Figure 4: Visualization of Prompt-CAM on different datasets. We show the top four attention maps (from left to right) per correctly classified test example triggered by the ground-truth classes.
@@ 3.1 Experimental Setup
Model. We consider three pre-trained ViT backbones, DINO [4], DINOv2 [29], and BioCLIP [38] across different scales including ViT-B (the main one we use) and ViT-S. The backbones are kept completely frozen when applying Prompt-CAM. We mainly used DINO, unless stated otherwise. More details can be found in Suppl.
@@ 3.2 Experiment Results
Is Prompt-CAM faithful? We first investigate whether Prompt-CAM highlights the image regions that the corresponding classifier focuses on when making predictions. We use Prompt-CAM to rank pixels based on the aggregated attention maps over the top heads. We then employ the insertion and deletion metrics [32], manipulating highly ranked pixels to measure confidence increase and drop.
For comparison, we consider post-hoc explainable methods like Grad-CAM [37], Eigen-CAM [25], Layer-CAM [13], and attention roll-out [14], based on the same ViT backbone with Linear Probing. As summarized in Table 1, Prompt-CAM yields higher insertion scores and lower deletion scores, indicating a stronger focus on discriminative image traits and highlighting Prompt-CAM’s enhanced interpretability over standard post-hoc algorithms.
@@
Grad-CAM [37] 0.52 0.17
Layer-CAM[13] 0.54 0.13
Eigen-CAM [25] 0.56 0.22
Attention roll-out [14] 0.55 0.27
@@
Linear Probing 99.2 78.6 82.4 92.4
Prompt-CAM 98.8 73.2 81.1 91.3
Prompt-CAM excels in trait identification (human assessment). We then conduct a quantitative human study to evaluate trait identification quality for Prompt-CAM, TesNet [47], and ProtoConcepts [21]. Participants with no prior knowledge about the algorithms were instructed to compare the expert-identified traits (in text, such as orange belly) and the top heatmaps generated by each method. If an expert-identified trait is seen in the heatmaps, it is considered identified by the algorithm. On average, participants recognized $60.49\%$ of traits for Prompt-CAM, significantly outperforming TesNet and ProtoConcepts whose recognition rates are $39.14\%$ and $30.39$%, respectively. The results highlight Prompt-CAM’s superiority in emphasizing and conveying relevant traits effectively. More details are in Suppl.
Classification accuracy comparison. We observe that Prompt-CAM shows a slight accuracy drop compared to Linear Probing (see Table 2). However, the images misclassified by Prompt-CAM but correctly classified by Linear Probing align with our design philosophy: Prompt-CAM classifies images based on the presence of class-specific, localized traits and would fail if they are invisible. As shown in Figure 5, discriminative traits—such as the red breast of the Red-breasted Grosbeak—are barely visible in images misclassified by Prompt-CAM due to occlusion, unusual poses, or lighting conditions. Linear Probing correctly classifies them by leveraging global information such as body shapes and backgrounds. Please see more analysis in Suppl.
Figure 5: Images misclassified by Prompt-CAM but correctly classified by Linear Probing. Species-specific traits—such as the red breast of “Red-breasted Grosbeak”—are barely visible in misclassified images while Linear Probing uses global features such as body shapes, poses, and backgrounds for correct predictions.
@@
Figure 6: Comparison of interpretable models. Visual demonstration (heatmaps and bounding boxes) of the four most activated responses of attention heads (Prompt-CAM and INTR) or prototypes of each method on a “Lazuli Bunting” example image.